Solve for x and y (2x/3) - (3y/4) = (11/12) and (x/3) + (7y/18) = (1/2)

Answer by mananth(16946)

(2x/3) - (3y/4) = (11/12)
LCM is 12
(4*2x-3*3y) /12 = 11/12
()8x-9y /12 = 11/12
multiply by 12
8x-9y=11..........................1
(x/3) + (7y/18) = (1/2)
LCM =18
(6*x+7y) /18 = 1/2
6x+7y= 18/2
6x+7y=9............................2
..
multiply equation 1 by 7 and equation 2 by 9 and add them to eliminate y
56x-63y+54x+63y=77+81
110x=158
x= 158/110
x=79/55
..
plug value of x in equation 2
6*79/55 +7y=9
474+385y=9*55
385y=495-474
385y=21
y= 21/385
y= 3/55